OpenStudy (just_gymnast_me):

Brian is creating a collage on a piece of cardboard that has an area of 160r4 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the seventh power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (anonymous):

T = (160)^2 (r^4)^2 / r^7

OpenStudy (anonymous):

@Niguyver please do not give direct answers. Let T represent the total number of pieces of paper used....T= 160r^4 / square root of r^7 ,,,,,Square the expression,to get rid of the square root,,, T = (160)^2 (r^4)^2 / r^7 ,,,,,, T = 25600r^8 / r^7 ,,,,, T = 25600r....Therefore,the expression of the total number of pieces of paper used is T = 25600r. total area:160r^4 area of one piece = r 7 − − so total pieces = 160r 4 r 7 − − now u just need to simplify... i will give a hint... a =a 1/2 and a x a y =a x+y okk... r 4 r 7 − − =r 4−7/2 =r 1/2 =r so final answer is 160r Answer is: 160r Hope this helps, have a wonderful day! If you need anymore math help please let me know, msg me, tag me, etc.

OpenStudy (just_gymnast_me):

thank you so much! I really appreciate it (:

OpenStudy (anonymous):

so I didn't help you

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I gave you the equation to find the total number of paper used

OpenStudy (just_gymnast_me):

You did help, thank you. (:

OpenStudy (anonymous):

cool